OpenStudy (anonymous):

the scale factor between a pair of similar triangles is one half. If the larger triangle has an area of 24 square units, what is the area of the smaller triangle?

OpenStudy (dumbcow):

The ratio of areas of similar triangles is always the squared ratio of their lengths (1/2)^2 = 1/4 24*(1/4) = 6